# A parametric finite element method for a degenerate multi-phase Stefan problem with triple junctions

### v1

@@ 2 Gradient Flow and Energy Dissipation Properties
where $\mathscr{H}^{d-1}$ denotes the $(d-1)$-dimensional Hausdorff measure.
@@
$\mathcal{M}_{m}$ :=$:=\left\{(\chi_{1},\cdots,\chi_{I_{R}})\in BV(\Omega;\{0,1\})^{I_{R}}\biggm{|}\sum_{\ell=1}^{I_{R}}\int_{\mathcal{R}_{\ell}[\Gamma(t)]}\beta_{\ell}\chi_{\ell}\,{\rm d}\mathscr{L}^{d}=m\right\},$
$T_{\Gamma}\mathcal{M}_{m}$ :=$:=\left\{V:\Gamma\to\mathbb{R}\biggm{|}\sum_{i=1}^{I_{S}}\int_{\Gamma_{i}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}V_{i}\,{\rm d}\mathscr{H}^{d-1}=0\right\},$
@@
$V_{1},V_{2}\right>_{T_{\Gamma}\mathcal{M}_{m}}:=\int_{\Omega}\nabla w_{1}\cdot\nabla w_{2}\,{\rm d}\mathscr{L}^{d}=-\sum_{i=1}^{I_{S}}\int_{\Gamma_{i}}w_{1}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}V_{2}\,{\rm d}\mathscr{H}^{d-1},$
@@ 3 Weak formulation
00 0 $=\sum_{i=1}^{I_{S}}\int_{\Gamma_{i}(t)}(-\left[\nabla w\right]_{b_{i}^{-}}^{b_{i}^{+}}\cdot\vec{\nu}_{i})\varphi\,{\rm d}\mathscr{H}^{d-1}-\int_{\Omega}\nabla w\cdot\nabla\varphi\,{\rm d}\mathscr{L}^{d}$
@@
$u,\,v\right>_{\Omega}:=\int_{\Omega}u\,v\,{\rm d}\mathscr{L}^{d}\quad\text{and}\quad\left<u,\,v\right>_{\Gamma(t)}:=\sum_{i=1}^{I_{S}}\int_{\Gamma_{i}(t)}u_{i}\,v_{i}\,{\rm d}\mathscr{H}^{d-1}.$
@@
$\nabla w,\,\nabla\varphi\right>_{\Omega}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<V_{i},\,\varphi\right>_{\Gamma_{i}(t)}=0.$ (3.3a)
@@
$\varkappa_{\varsigma}\vec{\nu}_{,}\,\vec{\eta}\right>_{\Gamma(t)}+\left<\varsigma\nabla_{\Gamma}\vec{\rm id},\,\nabla_{\Gamma}\vec{\eta}\right>_{\Gamma(t)}=0.$ (3.3c)
@@ 4 Parametric finite element method
We split the interval $[0,T]$ into $M$ sub-intervals $[t_{m-1},t_{m}]$ with $1\leq m\leq M$ whose lengths are equal to $\tau_{m}$. Then, we construct a fully discrete scheme which yields an approximate solution $(W^{m+1},\vec{X}^{m+1},\kappa^{m+1}_{\varsigma})$, provided that the previous discrete cluster $\Gamma^{m}$ is given, with $\Gamma^{m+1}=\vec{X}^{m+1}(\Gamma^{m})$. In order to describe $\Gamma^{m}$, and the discrete matching conditions that have to hold on the triple junction, we let $\Upsilon_{i}^{h}\,(i\in\mathbb{N}_{\leq I_{S}})$ be polyhedral reference surfaces with $\overline{\Upsilon_{i}^{h}}=\bigcup_{j=1}^{J_{i}}\overline{\sigma_{i,j}}$, where $\{\sigma_{i,j}\}_{j=1}^{J_{i}}$ is a family of mutually disjoint open $(d-1)$-simplices with vertices $\{\vec{q}_{i,k}\}_{k=1}^{K_{i}}$ vertices.
@@
$V^{h}(\Gamma^{m}_{i})$ :=$:=\left\{v\in C(\Gamma^{m}_{i})\biggm{|}v\mid_{\sigma^{m}_{i,j}}\mbox{ is affine}\quad\forall j\in\mathbb{N}_{\leq J_{i}}\right\},\qquad i\in\mathbb{N}_{\leq I_{S}},$
[…]
@@
$\underline{V}^{h}(\Gamma^{m})$ :=$:=\left\{(\vec{X}_{1},\cdots,\vec{X}_{I_{S}})\in\bigotimes_{i=1}^{I_{S}}\underline{V}^{h}(\Gamma^{m}_{i})\biggm{|}\vec{X}_{s^{k}_{1}}=\vec{X}_{s^{k}_{2}}=\vec{X}_{s^{k}_{3}}\quad\mbox{on}\quad\mathcal{T}_{k}^{m},\ \forall k\in\mathbb{N}_{\leq I_{T}}\right\},$
@@
$u,\,v\right>^{h}_{\Gamma^{m}_{i}}:=\frac{1}{d}\sum_{j=1}^{J_{i}}|\sigma^{m}_{i,j}|\sum_{k=1}^{d}\lim_{\begin{subarray}{c}\sigma^{m}_{i,j}\end{subarray}\ni\vec{q}\to\vec{q}^{m}_{i,j,k}}(uv)(\vec{q})\qquad i\in\mathbb{N}_{\leq I_{S}}.$
@@
$u,\,v\right>^{h}_{\Gamma^{m}}:=\sum_{i=1}^{I_{S}}\left<u_{i},\,v_{i}\right>^{h}_{\Gamma^{m}_{i}}.$ (4.3)
@@
$u,\,v\right>_{\Gamma^{m}}=\sum_{i=1}^{I_{S}}\left<u_{i},\,v_{i}\right>_{\Gamma^{m}_{i}}=\sum_{i=1}^{I_{S}}\int_{\Gamma^{m}_{i}}u_{i}v_{i}\,{\rm d}\mathscr{H}^{d-1}.$
@@
$\vec{\omega}^{m}_{i},\,\vec{\xi}\right>^{h}_{\Gamma^{m}_{i}}=\left<\vec{\nu}^{m}_{i},\,\vec{\xi}\right>_{\Gamma^{m}_{i}},\quad\forall\vec{\xi}\in\underline{V}^{h}(\Gamma^{m}_{i})\qquad i\in\mathbb{N}_{\leq I_{S}}.$
@@
$\nabla W^{m+1},\,\nabla\varphi\right>_{\Omega}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<\pi^{h}_{i}\left[\frac{\vec{X}^{m+1}_{i}-\vec{\rm id}}{\tau_{m}}\cdot\vec{\omega}^{m}_{i}\right],\,\varphi\right>^{(h)}_{\Gamma^{m}_{i}}=0.$ (4.4a)
@@
$\kappa^{m+1}_{\varsigma},\,\xi\right>^{h}_{\Gamma^{m}}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<W^{m+1},\,\xi\right>^{(h)}_{\Gamma^{m}_{i}}=0.$ (4.4b)
@@
$\kappa^{m+1}_{\varsigma}\vec{\omega}^{m},\,\vec{\eta}\right>^{h}_{\Gamma^{m}}+\left<\varsigma\nabla_{\Gamma}\vec{X}^{m+1},\,\nabla_{\Gamma}\vec{\eta}\right>_{\Gamma^{m}}=0,$ (4.4c)
@@
$\kappa_{\varsigma},\,\xi\right>^{h}_{\Gamma^{m}}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<W,\,\xi\right>^{(h)}_{\Gamma^{m}_{i}}=0\quad\forall\xi\in V^{h}(\Gamma^{m}),$ (4.5b)
$\kappa_{\varsigma}\vec{\omega}^{m},\,\vec{\eta}\right>^{h}_{\Gamma^{m}}+\left<\varsigma\nabla_{\Gamma}\vec{X},\,\nabla_{\Gamma}\vec{\eta}\right>_{\Gamma^{m}}=0\quad\forall\vec{\eta}\in\underline{V}^{h}(\Gamma^{m}).$ (4.5c)
@@
$\kappa_{\varsigma},\,\pi^{h}\left[\vec{X}\cdot\vec{\omega}^{m}\right]\right>^{h}_{\Gamma^{m}}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<W,\,\pi^{h}_{i}\left[\vec{X}\cdot\vec{\omega}^{m}_{i}\right]\right>^{(h)}_{\Gamma^{m}_{i}}$ $=0,$ (4.7)
[…]
@@
$\kappa^{m+1}_{\varsigma}\vec{\omega}^{m},\,\vec{X}^{m+1}-\vec{\rm id}\right>^{h}_{\Gamma^{m}}=-\left<\varsigma\nabla_{\Gamma}\vec{X}^{m+1},\,\nabla_{\Gamma}(\vec{X}^{m+1}-\vec{\rm id})\right>_{\Gamma^{m}}.$ (4.11)
@@
$\nabla_{\Gamma}\vec{X}^{m+1}_{i},\,\nabla_{\Gamma}(\vec{X}^{m+1}_{i}-\vec{\rm id})\right>_{\Gamma^{m}_{i}}\geq|\Gamma^{m+1}_{i}|-|\Gamma^{m}_{i}|,$ (4.12)
@@ 5 Matrix form of the linear system
$\begin{pmatrix}\tau_{m}A_{\Omega}&O&\vec{N}_{\Omega,\Gamma}^{T}\uuline{P}\\ B_{\Omega,\Gamma}&C_{\Gamma}&O\\ O&\uuline{P}\vec{D}_{\Gamma}&\uuline{P}\uuline{E_{\Gamma}}\uuline{P}\end{pmatrix}\begin{pmatrix}W^{m+1}\\ \kappa^{m+1}_{\varsigma}\\ \delta\vec{X}^{m+1}\end{pmatrix}=\begin{pmatrix}O\\ O\\ \mbox{\boldmath{$-$}}\uuline{P}\uuline{E_{\Gamma}}\uuline{P}\vec{X}^{m}\end{pmatrix},$ (5.1)
@@
$\vec{N}_{\Omega,\Gamma}:=\begin{pmatrix}\left[\beta\right]_{b_{1}^{-}}^{b_{1}^{+}}\vec{N}_{1}\\ \left[\beta\right]_{b_{2}^{-}}^{b_{2}^{+}}\vec{N}_{2}\\ \left[\beta\right]_{b_{3}^{-}}^{b_{3}^{+}}\vec{N}_{3}\\ \end{pmatrix},\quad$ $B_{\Omega,\Gamma}:=\begin{pmatrix}\left[\beta\right]_{b_{1}^{-}}^{b_{1}^{+}}B_{1}\\ \left[\beta\right]_{b_{2}^{-}}^{b_{2}^{+}}B_{2}\\ \left[\beta\right]_{b_{3}^{-}}^{b_{3}^{+}}B_{3}\\ \end{pmatrix},\quad C_{\Gamma}:=\begin{pmatrix}C_{1}&O&O\\ O&C_{2}&O\\ O&O&C_{3}\end{pmatrix},$
$\vec{D}_{\Gamma}:=\begin{pmatrix}\vec{D}_{1}&O&O\\ O&\vec{D}_{2}&O\\ O&O&\vec{D}_{3}\end{pmatrix},\quad$ $\uuline{E_{\Gamma}}:=\begin{pmatrix}\varsigma_{1}\uuline{E_{1}}&O&O\\ O&\varsigma_{2}\uuline{E_{2}}&O\\ O&O&\varsigma_{3}\uuline{E_{3}}\end{pmatrix}$
@@
$\left[B_{c}\right]_{k,j}:=\left<\Psi^{m}_{j},\,\Phi^{m}_{c,k}\right>^{(h)}_{\Gamma^{m}_{c}},\qquad$ $\left[C_{c}\right]_{k,\ell}:=\left<\Phi^{m}_{c,\ell},\,\Phi^{m}_{c,k}\right>^{(h)}_{\Gamma^{m}_{c}},$
$\left[\vec{D}_{c}\right]_{k,\ell}:=\left<\Phi^{m}_{c,\ell},\,\Phi^{m}_{c,k}\right>^{h}_{\Gamma^{m}_{c}}\vec{\omega}^{m}_{c,\ell},\qquad$ $\left[\uuline{E_{c}}\right]_{k,\ell}:=\left<\nabla_{\Gamma}\Phi^{m}_{c,\ell},\,\nabla_{\Gamma}\Phi^{m}_{c,k}\right>_{\Gamma^{m}_{c}}$

### v2

@@ 2 Gradient Flow and Energy Dissipation Properties
where $\mathscr{H}^{d-1}$ denotes the (d1)(d-1)-dimensional Hausdorff measure.
@@
$\mathcal{M}_{m}$ $:=\left\{(\chi_{1},\cdots,\chi_{I_{R}})\in BV(\Omega;\{0,1\})^{I_{R}}\biggm|\sum_{\ell=1}^{I_{R}}\int_{\mathcal{R}_{\ell}[\Gamma(t)]}\beta_{\ell}\chi_{\ell}\,{\rm d}\mathscr{L}^{d}=m\right\},$
$T_{\Gamma}\mathcal{M}_{m}$ $:=\left\{V:\Gamma\to\mathbb{R}\biggm|\sum_{i=1}^{I_{S}}\int_{\Gamma_{i}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}V_{i}\,{\rm d}\mathscr{H}^{d-1}=0\right\},$
@@
$\left<V_{1},V_{2}\right>_{T_{\Gamma}\mathcal{M}_{m}}:=\int_{\Omega}\nabla w_{1}\cdot\nabla w_{2}\,{\rm d}\mathscr{L}^{d}=-\sum_{i=1}^{I_{S}}\int_{\Gamma_{i}}w_{1}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}V_{2}\,{\rm d}\mathscr{H}^{d-1},$
@@ 3 Weak formulation
$0$ $=\sum_{i=1}^{I_{S}}\int_{\Gamma_{i}(t)}(-\left[\nabla w\right]_{b_{i}^{-}}^{b_{i}^{+}}\cdot\vec{\nu}_{i})\varphi\,{\rm d}\mathscr{H}^{d-1}-\int_{\Omega}\nabla w\cdot\nabla\varphi\,{\rm d}\mathscr{L}^{d}$
@@
$\left<u,\,v\right>_{\Omega}:=\int_{\Omega}u\,v\,{\rm d}\mathscr{L}^{d}\quad\text{and}\quad\left<u,\,v\right>_{\Gamma(t)}:=\sum_{i=1}^{I_{S}}\int_{\Gamma_{i}(t)}u_{i}\,v_{i}\,{\rm d}\mathscr{H}^{d-1}.$
@@
$\left<\nabla w,\,\nabla\varphi\right>_{\Omega}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<V_{i},\,\varphi\right>_{\Gamma_{i}(t)}=0.$ (3.3a)
@@
$\left<\varkappa_{\varsigma}\vec{\nu}_{,}\,\vec{\eta}\right>_{\Gamma(t)}+\left<\varsigma\nabla_{\Gamma}\vec{\rm id},\,\nabla_{\Gamma}\vec{\eta}\right>_{\Gamma(t)}=0.$ (3.3c)
@@ 4 Parametric finite element method
We split the interval $[0,T]$ into $M$ sub-intervals $t_{m-1},t_{m}]$ with $1\leq m\leq M$ whose lengths are equal to $\tau_{m}$. Then, we construct a fully discrete scheme which yields an approximate solution $(W^{m+1},\vec{X}^{m+1},\kappa^{m+1}_{\varsigma})$, provided that the previous discrete cluster $\Gamma^{m}$ is given, with $\Gamma^{m+1}=\vec{X}^{m+1}(\Gamma^{m})$. In order to describe $\Gamma^{m}$, and the discrete matching conditions that have to hold on the triple junction, we let $\Upsilon_{i}^{h}\,(i\in\mathbb{N}_{\leq I_{S}})$ be polyhedral reference surfaces with $\overline{\Upsilon_{i}^{h}}=\bigcup_{j=1}^{J_{i}}\overline{\sigma_{i,j}}$, where $\{\sigma_{i,j}\}_{j=1}^{J_{i}}$ is a family of mutually disjoint open (d1)(d-1)-simplices with vertices $\{\vec{q}_{i,k}\}_{k=1}^{K_{i}}$ vertices.
@@
$V^{h}(\Gamma^{m}_{i})$ $:=\left\{v\in C(\Gamma^{m}_{i})\biggm|v\mid_{\sigma^{m}_{i,j}}\mbox{ is affine}\quad\forall j\in\mathbb{N}_{\leq J_{i}}\right\},\qquad i\in\mathbb{N}_{\leq I_{S}},$
$\underline{V}^{h}(\Gamma^{m}_{i})$ $:=\left\{\mbox{{$v$}}=(v_{1},\cdots,v_{d})\in C(\Gamma^{m}_{i};\mathbb{R}^{d})\biggm|v_{k}\in V^{h}(\Gamma^{m}_{i})\quad\forall k\in\mathbb{N}_{\leq d}\right\},\qquad i\in\mathbb{N}_{\leq I_{S}}.$
@@
$\underline{V}^{h}(\Gamma^{m})$ $:=\left\{(\vec{X}_{1},\cdots,\vec{X}_{I_{S}})\in\bigotimes_{i=1}^{I_{S}}\underline{V}^{h}(\Gamma^{m}_{i})\biggm|\vec{X}_{s^{k}_{1}}=\vec{X}_{s^{k}_{2}}=\vec{X}_{s^{k}_{3}}\quad\mbox{on}\quad\mathcal{T}_{k}^{m},\ \forall k\in\mathbb{N}_{\leq I_{T}}\right\},$
@@
$\left<u,\,v\right>^{h}_{\Gamma^{m}_{i}}:=\frac{1}{d}\sum_{j=1}^{J_{i}}|\sigma^{m}_{i,j}|\sum_{k=1}^{d}\lim_{\begin{subarray}{c}\sigma^{m}_{i,j}\end{subarray}\ni\vec{q}\to\vec{q}^{m}_{i,j,k}}(uv)(\vec{q})\qquad i\in\mathbb{N}_{\leq I_{S}}.$
@@
$\left<u,\,v\right>^{h}_{\Gamma^{m}}:=\sum_{i=1}^{I_{S}}\left<u_{i},\,v_{i}\right>^{h}_{\Gamma^{m}_{i}}.$ (4.3)
@@
$\left<u,\,v\right>_{\Gamma^{m}}=\sum_{i=1}^{I_{S}}\left<u_{i},\,v_{i}\right>_{\Gamma^{m}_{i}}=\sum_{i=1}^{I_{S}}\int_{\Gamma^{m}_{i}}u_{i}v_{i}\,{\rm d}\mathscr{H}^{d-1}.$
@@
$\left<\vec{\omega}^{m}_{i},\,\vec{\xi}\right>^{h}_{\Gamma^{m}_{i}}=\left<\vec{\nu}^{m}_{i},\,\vec{\xi}\right>_{\Gamma^{m}_{i}},\quad\forall\vec{\xi}\in\underline{V}^{h}(\Gamma^{m}_{i})\qquad i\in\mathbb{N}_{\leq I_{S}}.$
@@
$\left<\nabla W^{m+1},\,\nabla\varphi\right>_{\Omega}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<\pi^{h}_{i}\left[\frac{\vec{X}^{m+1}_{i}-\vec{\rm id}}{\tau_{m}}\cdot\vec{\omega}^{m}_{i}\right],\,\varphi\right>^{(h)}_{\Gamma^{m}_{i}}=0.$ (4.4a)
@@
$\left<\kappa^{m+1}_{\varsigma},\,\xi\right>^{h}_{\Gamma^{m}}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<W^{m+1},\,\xi\right>^{(h)}_{\Gamma^{m}_{i}}=0.$ (4.4b)
@@
$\left<\kappa^{m+1}_{\varsigma}\vec{\omega}^{m},\,\vec{\eta}\right>^{h}_{\Gamma^{m}}+\left<\varsigma\nabla_{\Gamma}\vec{X}^{m+1},\,\nabla_{\Gamma}\vec{\eta}\right>_{\Gamma^{m}}=0,$ (4.4c)
@@
$\left<\kappa_{\varsigma},\,\xi\right>^{h}_{\Gamma^{m}}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<W,\,\xi\right>^{(h)}_{\Gamma^{m}_{i}}=0\quad\forall\xi\in V^{h}(\Gamma^{m}),$ (4.5b)
$\left<\kappa_{\varsigma}\vec{\omega}^{m},\,\vec{\eta}\right>^{h}_{\Gamma^{m}}+\left<\varsigma\nabla_{\Gamma}\vec{X},\,\nabla_{\Gamma}\vec{\eta}\right>_{\Gamma^{m}}=0\quad\forall\vec{\eta}\in\underline{V}^{h}(\Gamma^{m}).$ (4.5c)
@@
$\left<\kappa_{\varsigma},\,\pi^{h}\left[\vec{X}\cdot\vec{\omega}^{m}\right]\right>^{h}_{\Gamma^{m}}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<W,\,\pi^{h}_{i}\left[\vec{X}\cdot\vec{\omega}^{m}_{i}\right]\right>^{(h)}_{\Gamma^{m}_{i}}$ $=0,$ (4.7)
$\left<\kappa_{\varsigma}\vec{\omega}^{m},\,\vec{X}\right>^{h}_{\Gamma^{m}}+\left<\varsigma\nabla_{\Gamma}\vec{X},\,\nabla_{\Gamma}\vec{X}\right>_{\Gamma^{m}}=0.$ (4.8)
@@
$\left<\kappa^{m+1}_{\varsigma}\vec{\omega}^{m},\,\vec{X}^{m+1}-\vec{\rm id}\right>^{h}_{\Gamma^{m}}=-\left<\varsigma\nabla_{\Gamma}\vec{X}^{m+1},\,\nabla_{\Gamma}(\vec{X}^{m+1}-\vec{\rm id})\right>_{\Gamma^{m}}.$ (4.11)
@@
$\left<\nabla_{\Gamma}\vec{X}^{m+1}_{i},\,\nabla_{\Gamma}(\vec{X}^{m+1}_{i}-\vec{\rm id})\right>_{\Gamma^{m}_{i}}\geq|\Gamma^{m+1}_{i}|-|\Gamma^{m}_{i}|,$ (4.12)
@@ 5 Matrix form of the linear system
$\begin{pmatrix}\tau_{m}A_{\Omega}&O&\vec{N}_{\Omega,\Gamma}^{T}\uuline{P}\\ B_{\Omega,\Gamma}&C_{\Gamma}&O\\ O&\uuline{P}\vec{D}_{\Gamma}&\uuline{P}\uuline{E_{\Gamma}}\uuline{P}\end{pmatrix}\begin{pmatrix}W^{m+1}\\ \kappa^{m+1}_{\varsigma}\\ \delta\vec{X}^{m+1}\end{pmatrix}=\begin{pmatrix}O\\ O\\ \mbox{{$-$}}\uuline{P}\uuline{E_{\Gamma}}\uuline{P}\vec{X}^{m}\end{pmatrix}$, (5.1)
@@
$\vec{N}_{\Omega,\Gamma}:=\begin{pmatrix}\left[\beta\right]_{b_{1}^{-}}^{b_{1}^{+}}\vec{N}_{1}\\ \left[\beta\right]_{b_{2}^{-}}^{b_{2}^{+}}\vec{N}_{2}\\ \left[\beta\right]_{b_{3}^{-}}^{b_{3}^{+}}\vec{N}_{3}\\ \end{pmatrix}$$\quad$ $B_{\Omega,\Gamma}:=\begin{pmatrix}\left[\beta\right]_{b_{1}^{-}}^{b_{1}^{+}}B_{1}\\ \left[\beta\right]_{b_{2}^{-}}^{b_{2}^{+}}B_{2}\\ \left[\beta\right]_{b_{3}^{-}}^{b_{3}^{+}}B_{3}\\ \end{pmatrix},\quad C_{\Gamma}:=\begin{pmatrix}C_{1}&O&O\\ O&C_{2}&O\\ O&O&C_{3}\end{pmatrix}$,
$\vec{D}_{\Gamma}:=\begin{pmatrix}\vec{D}_{1}&O&O\\ O&\vec{D}_{2}&O\\ O&O&\vec{D}_{3}\end{pmatrix}$$\quad$ $\uuline{E_{\Gamma}}:=\begin{pmatrix}\varsigma_{1}\uuline{E_{1}}&O&O\\ O&\varsigma_{2}\uuline{E_{2}}&O\\ O&O&\varsigma_{3}\uuline{E_{3}}\end{pmatrix}$
@@
$\left[B_{c}\right]_{k,j}:=\left<\Psi^{m}_{j},\,\Phi^{m}_{c,k}\right>^{(h)}_{\Gamma^{m}_{c}}$$\qquad$ $\left[C_{c}\right]_{k,\ell}:=\left<\Phi^{m}_{c,\ell},\,\Phi^{m}_{c,k}\right>^{(h)}_{\Gamma^{m}_{c}}$,
$\left[\vec{D}_{c}\right]_{k,\ell}:=\left<\Phi^{m}_{c,\ell},\,\Phi^{m}_{c,k}\right>^{h}_{\Gamma^{m}_{c}}\vec{\omega}^{m}_{c,\ell}$$\qquad$ $\left[\uuline{E_{c}}\right]_{k,\ell}:=\left<\nabla_{\Gamma}\Phi^{m}_{c,\ell},\,\nabla_{\Gamma}\Phi^{m}_{c,k}\right>_{\Gamma^{m}_{c}}$